a population of pwd learners in 5 schools in a certain municipality of quezon province are (x) 9 5 6 12 and 25. suppose that two schools were selected as sample determine the mean and variance of the sampling disteibution of sample mean

1
Expert's answer
2022-03-20T06:44:39-0400

"EX=\frac{9+5+6+12+25}{5}=11.4\\EX^2=\frac{9^2+5^2+6^2+12^2+25^2}{5}=182.2\\DX=EX^2-\left( EX \right) ^2=182.2-11.4^2=52.24\\\bar{x}=EX=11.4\\s^2=\frac{DX}{2}=\frac{52.24}{2}=26.12\\"